A basketball player makes a jump shot. The 0.650 kg ball is released at a height of 2.10 m above the floor with a speed of 7.40 m/s. The ball goes through the net 3.10 m above the floor at a speed of 4.20 m/s. What is the work done on the ball by air resistance, a nonconservative force
Answer #1

intial energy = KE + PE

= 0.5mv^2 + mgh

= 0.65 ( 0.5*7.4^2 + 9.8*2.1 )

= 31.174 J

final Energy = 0.65*(0.5*4.2^2 + 9.8*3.1)

= 25.48 J

this loss is energy is due to work done on the ball by air resistance

so W = (31.174 - 25.48) = 5.694 J